.Two identical ball-head on. The initial velocity of one is 0.75m/s ⁄ –East , while that of the other one is 0.43m/s ⁄ -West. If the collision is perfectly elastic. Determine the velocity of each ball.
1
Expert's answer
2020-07-08T13:14:20-0400

According to the law of conservation of momentum (in projection on the x-axis directed to the East):


"mv_1 -mv_2 = mv_1' + mv_2'\\
v_1 - v_2 = v_1' + v_2'"

where "v_1" and "v_2" are velocities of the first and second ball respectively before the collision, "v_1'" and "v_2'" - after the collision. Since masses are identical they can be canceled.

Since the collision is perfectly elastic, the energy conservation law holds:


"\dfrac{mv_1^2}{2} + \dfrac{mv_2^2}{2} = \dfrac{mv_1'^2}{2} + \dfrac{mv_2'^2}{2}\\
v_1^2 + v_2^2 = v_1'^2 + v_2'^2"

Thus, have two equations for two variables: "v_1'" and "v_2'".


"v_1 - v_2 = v_1' + v_2'\\
v_1^2 + v_2^2 = v_1'^2 + v_2'^2"

Substituting the values, obtain:


"v_1' + v_2' = 0.75-0.43=0.32\\
v_1^2 + v_2^2 = 0.75^2 + 0.43^2=0.7474"

Expressing "v_1'" from the first equation and substituting it to the second one, get:


"v_1' = 0.32-v_2'\\
( 0.32-v_2')^2 + v_2'^2 = 0.7474\\
0.1024-0.64v_2' + 2v_2'^2 = 0.7474\\
 2v_2'^2-0.64v_2'-0.645 = 0"


The solutions are:


"v_2' =- 0.43m/s \\
v_2' =0.75m/s"

We should chose the last one, for the ball cannot not to change its velocity after the collision.

Then:


"v_1' = 0.32-0.75 = -0.43m\/s"

Answer. The velocities after the collision: 1st ball 0.43 m/s to the West, 2nd ball 0.75 m/s to the East.